OpenStudy (anonymous):

How do I solve this? a (x-3)+8 = b (x-1)

OpenStudy (aric200):

1) a(x-3)+8=b(x-1) ax-3a+8 = bx-b ax-bx = 3a-b-8 x(a-b)= 3a-b-8 x = (3a-b-8)/(a-b)

OpenStudy (anonymous):

a x-3 a+8=b x-b a x-b x=3 a-b-8 ?